# Catch Me if You Can: Detecting Unauthorized Data Use in Deep Learning Models

### v1

@@ 4.4.2 Set-based membership inference for reliable data provenance
4.4.3 Provenance verification process
@@
The first step is to derive an inference threshold $c$ based on a pre-defined FPR ($\alpha$). $\alpha$ is usually a small number such as 0.1% or 0% to ensure low false positive (e.g., Section 5 shows that MembershipTracker can support the users to achieve 100% TPR with 0% FPR in many cases). To derive the threshold $c$, $U$ first computes the empirical loss histogram for data in the out world $\mathcal{P}_{\text{out}}$. For each non-member user in $\mathcal{P}_{\text{out}}$, the auditing user ${U}$ first computes the per-instance loss on their data (line 2), and then derives the average loss for each (line 3). Line 4 estimates the CDF for the loss distribution, and computes its $\alpha$-percentile to derive a threshold $c$ [70].
@@
$\alpha$: FPR to control the verification process (e.g., 0.1%);
@@
5 Evaluation
Datasets and model training. We consider six common benchmark datasets, including CIFAR100 [35], CIFAR10 [35], ArtBench [41], CelebA [45], TinyImageNet [36] and ImageNet-1k [19]. This section considers the first five datasets and the evaluation on the ImageNet-1k dataset is in Section 6. We follow standard model training practice such as using data augmentations and validation set - the model training details are in Appendix A.
@@ 5 Evaluation
As explained earlier, we create outlier features as random color stripes, each with 16 stripes from 11 different common colors. Such a large feature set ($11^{16}$ choices) can support multiple users to generate their own features for data marking, and we use a image blending ratio $m=0.7$. Next, we follow Co et al. [17] to generate random perlin noise for each sample with $\delta=8/255$. We provide additional visualizations of the original and marked samples in Appendix A.
@@
To perform MI, we directly use the cross-entropy loss as the signal function value, which is easy to compute by the users. We consider the global-threshold method by Carlini et al. [8] as the baseline method; and we do not choose the shadow-model-based approaches [8, 6, 70]. The reason is that, even though these methods can achieve better results, they still yield limited increase of MI success (Section 4.1); and more importantly, training shadow models is prohibitively challenging for ordinary personal users to carry out (Section 3.1.1).
The closest work to ours in improving the MI success without shadow-model calibration is Tramer et al. [65]. However, their technique assumes the users can manipulate the data labels, which excludes the real-world scenarios where the data labels are assigned by the model creator or an external service [21, 68, 19]. Though MembershipTracker does not impose such an assumption, for completeness, we still quantitatively compare it with the method by Tramer et al. (Appendix E).
5.1 Results
@@
Table 1: Evaluation on different datasets (parenthesized numbers show accuracy diff. compared with the original models).
@@
As shown, with MembershipTracker, the users can reliably trace the provenance of their marked data with 0% FPR@100% TPR. In comparison, without MembershipTracker, performing standard instance-based MI on the unmodified target samples incurs much higher FPR, with an average of 47.52% FPR@100% TPR. This is because: 1) the unmarked samples cannot be strongly memorized by the model; and 2) the instance-based MI process fails to leverage the collective information across the target samples by each user. These two limitations are overcome by MembershipTracker’s data marking and set-based MI process; and the detailed ablation study on these two components is presented in Appendix C.
@@
Evaluation on different architectures and training-set sizes. We next consider additional evaluation (using CIFAR100 dataset) on different model architectures (including WideResNet [72], ResNet [23], ResNext [69], SeNet [28], GoogleNet [63] and DenseNet [29]), as well as different training-set sizes (from 5,000 to 25,000). We observe a similar trend as before (thus we omit the discussion), and the detailed results are in Appendix B.
@@
Table 2: Evaluation on multiple target users. Supporting more target users indicates more specially-marked samples in the training set, which can lead to higher accuracy drop; but MembershipTracker still maintains high auditing performance.
@@
Supporting multiple users requires the model to memorize more target samples (from different users). However, DL models are known to be capable of memorizing a large corpus of data (e.g., to encode sensitive information [57, 14]). Likewise, in our case, the model still strongly memorizes the marked samples and enables MembershipTracker to achieve 0% FPR@100% TPR even under multiple target users.
Meanwhile, as the number of target users increases, the model also has higher accuracy drop. For instance, supporting 100 target users (i.e., 10% of training data are marked) incurs a 1.39% accuracy drop, while supporting 500 users has a higher 8.93% drop. This is understandable as supporting more users implies more training samples are marked.
5.2 Robustness to Countermeasures
@@
Table 3: Overview of the evaluated countermeasures.
@@
5.2.1 MI defense
DP-based defense. Differentially private (DP) training [3] is a principled defense that can bound the influence of any training samples to the model, by means of gradient clipping and noise injection to the clipped gradients. We adopt the implementation from Aerni et al. to consider the strong DP-SGD baselines that comprise various DP-training techniques such as custom initialization scheme, augmentation multiplicity [4]. We also follow their work to tune hyperparameters for high model utility. We use a tight clipping norm of 1, and inject different amounts of noise ($\sigma$) in the evaluation.
The results are in Table 4. DPSGD with a loose DP bound is known to be able to reduce the exposure from the instance-based MI methods that aim to de-identify some individual training instances [4, 8], and our results also validate this (e.g., in our context, the FPR@100% TPR under instance-based MI is increased to 47.22%, with $\sigma=0.01$). With that said, MembershipTracker can still overcome the defense via its set-based MI process, because the average loss of the samples by the member users, though increased due to defense, are still lower than that by the non-member users (average 0.1642 vs. 4.130). This avails MembershipTracker to maintain a low 0.01% FPR@100% TPR (and the defense already incurs 6.24% accuracy drop). Increasing the amount of noise can further increase the FPR by MembershipTracker, but it also inflicts a larger accuracy drop (undesirable).
Table 4: Evaluating MembershipTracker against DP-SGD with different amounts of noise injected.
@@ 5.2.1 MI defense
Empirical defense. Next, we also evaluate two representative defenses that aim to provide strong empirical privacy while preserving model utility: SELENA (USENIX’22) [64] and HAMP (NDSS’24) [15]. We follow the original work to set up both techniques. They both can mitigate the model’s memorization (increase the prediction loss) on the marked samples, and with small accuracy drop (2.17% and 1.42%). However, we find that the average loss of the samples by the member users are still much lower than that by the non-member users, and thus MembershipTracker still achieves a low FPR of 0.97% (on SELENA) and 0.83% (on HAMP).
@@
5.2.2 Adversarial augmentation
We consider employing additional augmentation techniques during training to mitigate the model’s memorization on the marked samples. We study two strategies, one by injecting Gaussian noise and the other by performing multiple random data augmentations (RandAugment [18]). For the former, we inject zero-mean noise with various standard deviation values (0.02$\sim$0.5); For RandAugment [18], we consider different number of augmentations (3$\sim$15).
Figure 6: Adv. augmentation by Gaussian noise injection and RandAugment [18]. MembershipTracker is resilient to moderate input augmentation, and aggressive augmentation can increase the TPR, but also jeopardize model accuracy.
@@ 5.2.2 Adversarial augmentation
5.2.3 Model fine-tuning
We now evaluate how model fine-tuning can mitigate the model’s memorization on the target samples. We assume an ideal defender with access to a small set (equivalent to 10% of the training set) of clean and unmodified samples. We fine-tune the model with various fine-tuning rates from 0.1$\sim$0.00001, with 20 epochs for each. The results are shown on the left of Fig. 7. When the fine-tuning rates are small ($<$0.05), the models can preserve high model accuracy, yet they still exhibit strong memorization on the target samples, and MembershipTracker still achieves 0% FPR@100% TPR.
@@ 5.2.3 Model fine-tuning
Figure 7: Defense evaluation on model fine-tuning and model pruning. MembershipTracker is resilient to moderate model fine-tuning and pruning. Aggressive countermeasures can slightly increase the FPR incurred by MembershipTracker, but at the cost of major accuracy degradation.
5.2.4 Model pruning
Another related countermeasure we consider is to prune the model’s parameters, with the goal of removing the parameters that are associated with the target samples. We use the popular pruning strategy by Han et al. [22], which sets the parameters with small absolute values to zero and has minimal impact to the model’s performance. We evaluate different pruning ratios and the results are on the right of Fig. 7.
We find that MembershipTracker is highly resilient even when 70% of the parameters are pruned, where MembershipTracker has only 0.02% FPR@100% TPR, but the defense already incurs a 11.8% accuracy drop. More aggressive pruning with a 80% pruning ratio can slightly increase the FPR to 1.16%, but with a much higher 36.9% accuracy drop.
5.2.5 Model-level defenses under multi-target setting
@@
Among different defenses, we find that model fine-tuning stands out as the most potent solution that outperforms other countermeasures in terms of high mitigation performance with low accuracy drop.
Compared with other methods (such as specialized MI defenses [3, 64, 15]), fine-tuning only needs to perform standard fine-tuning on a small set of samples, and it can increase the FPR@100% TPR to 21.4% (with a fine-tuning rate of 0.05) with a mere 2.09% accuracy drop.
@@
Figure 8: Unsupervised OOD detection. How different data-marking intensities affect the detection performance. Left: Samples marked with higher marking intensity are prone to be detected (0.8085 AUC). Right: Reducing the marking intensity can effectively reduce the AUC to only 0.6333.
5.2.6 Input detection
@@
Unsupervised OOD detection. We consider eight common detection techniques: Mahalanobis [37], MSP [25], Energy-based OOD [42], RMD [51], kNN [62], KLMatching [24], VIM [67], DICE [61]. We follow the implementation from [34], and use a pre-trained ImageNet model to produce the feature representation. As before, we assume an ideal defender with access to a small set of clean and unmodified samples (equivalent to 10% of the training set) for fitting the unsupervised detector, which does not assume any prior knowledge of the potential outlier samples.
We report the results on the left of Fig. 8, where, for brevity, we report only the highest AUC among all detectors. We find that the Mahalanobis-based approach [37] achieves the highest performance with 0.8085 AUC. In light of this, we next explain how to (partially) mitigate it.
@@
Naturally, reducing the marking intensity would attenuate the model’s memorization on the marked sample, and increase their prediction loss. But we find that the prediction loss on the member samples are still much lower than that on the non-members, and thus MembershipTracker can overcome it via the set-based MI process (similar to Section 5.2.1), and maintain a 0% FPR@100% TPR [^2] However, there are certain scenarios where the reduced data-marking intensity would lead to lower auditing performance, and an evaluation for this is in Appendix C.1.2..
Figure 9: Supervised OOD detection. Blue line assumes a defender with perfect knowledge of the type of outlier features adopted by the users for data marking. Orange line considers a realistic defender with knowledge of the possible types of outlier features, but the exact one followed by the users to mark their data is unpredictable.
Supervised OOD detection. Next, we consider a more knowledgeable defender who is aware of the potential data marking strategy by MembershipTracker. We use the MCHAD method to train a supervised OOD detector [33], and we consider two separate scenarios below.
@@
Specifically, we train the detector on samples marked with random color stripes, and evaluate it on samples marked with the features from an OOD (TinyImageNet) dataset (this is a similarly effective alternative strategy we study in Appendix C.1.3). The result is in the orange line in Fig. 9, and the detection performance degrades considerably under such a realistic setting (with only 0.66 AUC).
@@
Spurious features detection is another related method. We evaluate a leading technique by Neuhaus et al. [49] based on neural PCA components. But we find that it also has limited success in detecting the marked samples (0.6181 AUC).
5.2.7 Output perturbation
MembershipTracker leverages the model’s outputs to perform MI auditing, and we consider two strategies to perturb the model’s outputs: (1) injecting Gaussian noise to the outputs; and (2) returning only the top-1 prediction label [16]. However, we find that MembershipTracker still maintains high auditing success even with the perturbed outputs.
For instance, aggressively injecting Gaussian noise (with zero mean and standard deviation $\sigma=3$) degrades the accuracy from 65.32% to 52.05%, under which MembershipTracker still maintains 0% FPR@100% TPR. Completely shielding the output probabilities is ineffective either, and MembershipTracker can still achieve 0.04% FPR@100% TPR in the label-only setting. The detailed evaluation can be referred to Appendix G.
6 Evaluation on ImageNet Training
This section evaluates MembershipTracker under the full-sized ImageNet-1k training (over 1.28 million training samples). We train a ResNet50 [23] and Swin-Transformer [44], and they have 75.09% and 80.29% accuracy respectively (both with <1% accuracy drop). We consider the challenging setting of supporting multiple users in different classes (1,000 target users in total). Each user contributes 125 samples to the training set, which is $<$0.01% in proportion.
We find that even under the large-scale training setting, the marked samples can still be strongly memorized by the models, which enables MembershipTracker to achieve 0% FPR@100% TPR for all the users.

### v2

@@ 4.4.2. Set-based membership inference for reliable data provenance
4.4.3. Provenance verification process
@@
The first step is to derive an inference threshold $c$ based on a pre-defined FPR ($\alpha$). $\alpha$ is a small number such as 0.1% or 0% to ensure low false positive (e.g., Section 5 shows that MembershipTracker can support the users to achieve 100% TPR with 0% FPR in many cases). To derive the threshold $c$, $U$ first computes the empirical loss histogram for data in the out world $\mathcal{P}_{\text{out}}$. For each non-member user in $\mathcal{P}_{\text{out}}$, the auditing user ${U}$ first computes the per-instance loss on their data (line 2), and then derives the average loss for each (line 3). Line 4 estimates the CDF for the loss distribution, and computes its $\alpha$-percentile to derive a threshold $c$ (Ye et al., 2022).
@@
$\alpha$: False positive rate (FPR) to control the verification process (e.g., 0.1%);
@@
5. Evaluation
Datasets and model training. We consider six common benchmark datasets, including CIFAR100 (Krizhevsky et al., 2009), CIFAR10 (Krizhevsky et al., 2009), ArtBench (Liao et al., 2022), CelebA (Liu et al., 2015), TinyImageNet (Le and Yang, 2015) and ImageNet-1k (Deng et al., 2009). This section considers the first five datasets and the evaluation on the ImageNet-1k dataset is in Section 6. We follow standard model training practice such as using data augmentations and validation set - the model training details are in Appendix A.
@@ 5. Evaluation
As explained earlier, we create outlier features as random color stripes, each with 16 stripes from 11 different common colors. Such a large feature set ($11^{16}$ choices) can support multiple users to generate their own features for data marking, and we use a image blending ratio $m=0.7$. Next, we follow Co et al. (Co et al., 2019) to generate random perlin noise for each sample with $\delta=8/255$. We provide additional visualizations of the original and marked samples in Appendix A.
@@
To perform MI, we directly use the cross-entropy loss as the signal function value, which is easy to compute by the users. We consider the global-threshold method by Carlini et al. (Carlini et al., 2022) as the baseline method; and we do not choose the shadow-model-based approaches (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022). The reason is that, even though these methods can achieve better results, they still yield limited increase of MI success (Section 4.1); and more importantly, training shadow models is prohibitively challenging for ordinary personal users to carry out (Section 3.1.1).
The closest work to ours in improving the MI success without shadow-model calibration is Tramer et al. (Tramèr et al., 2022). However, their technique assumes the users can manipulate the data labels, which excludes the real-world scenarios where the data labels are assigned by the model creator or an external service (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). Though MembershipTracker does not impose such an assumption, for completeness, we still quantitatively compare it with the method by Tramer et al. (Appendix E).
5.1. Results
@@
Table 1. Evaluation on different datasets (parenthesized numbers show accuracy diff. with the original models).
@@
As shown, with MembershipTracker, the users can reliably trace the provenance of their marked data with 0% FPR@100% TPR. In comparison, without MembershipTracker, performing standard instance-based MI on the unmodified target samples incurs much higher FPR, with an average of 47.52% FPR@100% TPR (we also validate that it similarly has low performance under the low FPR regime, e.g., $<1\%$ TPR@0.1% FPR). This is because: 1) the unmarked samples cannot be strongly memorized by the model; and 2) the instance-based MI process fails to leverage the collective information across the target samples by each user. These two limitations are overcome by MembershipTracker’s data marking and set-based MI process; and the detailed ablation study on these two components is presented in Appendix C.
@@
Evaluation on different architectures and training-set sizes. We next consider additional evaluation (using CIFAR100 dataset) on different model architectures (including WideResNet (Zagoruyko and Komodakis, 2016), ResNet (He et al., 2016), ResNext (Xie et al., 2017), SeNet (Hu et al., 2018), GoogleNet (Szegedy et al., 2015) and DenseNet (Huang et al., 2017)), as well as different training-set sizes (from 5,000 to 25,000). We observe a similar trend as before (thus we omit the discussion), and the detailed results are in Appendix B.
@@
Table 2. Evaluation on multiple target users. Supporting more users indicates more specially-marked samples in the dataset, which can lead to higher accuracy drop; but MembershipTracker still maintains high auditing performance.
@@
Supporting multiple users requires the model to memorize more target samples (from different users). However, DL models are known to be capable of memorizing a large corpus of data (e.g., to encode sensitive information (Song et al., 2017; Chen and Pattabiraman, 2024a)). Likewise, in our case, the model still strongly memorizes the marked samples and enables MembershipTracker to achieve 0% FPR@100% TPR even under multiple target users.
Meanwhile, as the number of target users increases, the model also has higher accuracy drop. E.g., supporting 100 target users (i.e., 10% of training set are marked) incurs a 1.39% accuracy drop, while supporting 500 users has a higher 8.93% drop. This is understandable as supporting more users implies more training samples are marked. Supporting more target users will result in a larger accuracy drop - we defer this to future work.
5.2. Robustness to Countermeasures
@@
Table 3. Overview of evaluated countermeasures.
@@
Table 4. Evaluating MembershipTracker against DP-SGD with different amounts of noise injected.
@@
5.2.1. MI defense
DP-based defense. Differentially private (DP) training (Abadi et al., 2016) is a principled defense that can bound the influence of any training samples to the model, by means of gradient clipping and noise injection to the clipped gradients. We adopt the implementation from Aerni et al. to consider the strong DP-SGD baselines that comprise various DP-training techniques such as custom initialization scheme, augmentation multiplicity (Aerni et al., 2024). We also follow their work to tune hyperparameters for high model utility. We use a tight clipping norm of 1, and inject different amounts of noise ($\sigma$).
The results are in Table 4. DPSGD with a loose DP bound is known to be able to reduce the exposure from the instance-based MI methods that aim to de-identify some individual training instances (Aerni et al., 2024; Carlini et al., 2022), and our results also validate this (e.g., in our context, the FPR@100% TPR under instance-based MI is increased to 47.22%, with $\sigma=0.01$). With that said, MembershipTracker can still overcome the defense via its set-based MI process, because the average loss of the samples by the member users, though increased due to defense, are still lower than that by the non-member users (average 0.1642 vs. 4.130). This avails MembershipTracker to maintain a low 0.01% FPR@100% TPR (and the defense already incurs 6.24% accuracy drop). Increasing the amount of noise can further increase the FPR by MembershipTracker, but it also inflicts a larger accuracy drop (undesirable).
Empirical defense. Next, we also evaluate two representative defenses that aim to provide strong empirical privacy while preserving model utility: SELENA (Tang et al., 2022) and HAMP (Chen and Pattabiraman, 2024b). We follow the original work to set up both techniques. They both can mitigate the model’s memorization (increase the prediction loss) on the marked samples, and with small accuracy drop (2.17% and 1.42%). However, we find that the average loss of the samples by the member users are still much lower than that by the non-member users, and thus MembershipTracker still achieves a low FPR of 0.97% (on SELENA) and 0.83% (on HAMP).
@@ 5.2.1. MI defense
5.2.2. Adversarial augmentation
We consider employing additional augmentation techniques during training to mitigate the model’s memorization on the marked samples. We study three strategies: (1) injecting Gaussian noise (we inject zero-mean noise with various standard deviation values (0.02$\sim$0.5)); (2) using Mixup (Zhang et al., 2018) (we use seven different alpha parameters from 0.1 to 1.5); and (3) using RandAugment (Cubuk et al., 2020) to perform multiple random data augmentations (we consider 3$\sim$15 different augmentations).
We find that MembershipTracker still maintains 0% FPR@100% TPR under Mixup training with different parameters. We report the results for the other two strategies in Fig. 6.
Figure 6. Adv. augmentation by Gaussian noise injection and RandAugment (Cubuk et al., 2020). MembershipTracker is resilient to moderate input augmentation, and aggressive augmentation can increase the TPR, but also jeopardize model accuracy.
@@ 5.2.2. Adversarial augmentation
5.2.3. Model fine-tuning
We now evaluate how model fine-tuning can mitigate the model’s memorization on the target samples. We assume an ideal defender with access to a small set (equivalent to 10% of training set) of clean and unmodified samples. We fine-tune the model with various fine-tuning rates from 0.1$\sim$0.00001, with 20 epochs for each. The results are on the left of Fig. 7. When the fine-tuning rates are small ($<$0.05), the models still exhibit strong memorization on the target samples, and thus MembershipTracker maintains 0% FPR@100% TPR.
@@ 5.2.3. Model fine-tuning
Figure 7. Defense evaluation on model fine-tuning and model pruning. MembershipTracker is resilient to moderate model fine-tuning and pruning. Aggressive countermeasures can slightly increase the FPR incurred by MembershipTracker, but at the cost of major accuracy degradation.
5.2.4. Model pruning
Another related countermeasure we consider is to prune the model’s parameters, with the goal of removing the parameters that are associated with the target samples. We use the popular pruning strategy by Han et al. (Han et al., 2015), which sets the parameters with small absolute values to zero and has minimal impact to the model’s performance. We evaluate different pruning ratios and the results are on the right of Fig. 7.
We find that MembershipTracker is highly resilient even when 70% of the parameters are pruned, where MembershipTracker has only 0.02% FPR@100% TPR, but the defense already incurs a 11.8% accuracy drop. More aggressive pruning can slightly increase the FPR, but with a much higher accuracy drop.
5.2.5. Model-level defenses under multi-target setting
@@
Among different defenses, we find that model fine-tuning stands out as the most potent solution that outperforms other methods in terms of high mitigation performance with low accuracy drop.
Compared with other methods (e.g., MI defenses (Abadi et al., 2016; Tang et al., 2022; Chen and Pattabiraman, 2024b)), fine-tuning only needs to perform standard fine-tuning on a small set of samples, and it can increase the FPR@100% TPR to 21.4% (with a fine-tuning rate of 0.05) with a mere 2.09% accuracy drop.
@@
5.2.6. Input detection
@@
Unsupervised OOD detection. We consider eight common detection techniques: Mahalanobis (Lee et al., 2018), MSP (Hendrycks and Gimpel, 2016), Energy-based OOD (Liu et al., 2020), RMD (Ren et al., 2021), kNN (Sun et al., 2022), KLMatching (Hendrycks et al., 2022), VIM (Wang et al., 2022), DICE (Sun and Li, 2022). We follow the implementation from (Kirchheim et al., 2022b), and use a pre-trained ImageNet model to produce the feature representation. As before, we assume an ideal defender with access to a small set of clean and unmodified samples (equivalent to 10% of the training set) for fitting the unsupervised detector, which does not assume any prior knowledge of the potential outlier samples.
We report the results on the left of Fig. 8, where, for brevity, we report only the highest AUC among all detectors. We find that the Mahalanobis-based approach (Lee et al., 2018) achieves the highest performance with 0.8085 AUC. Next, we explain how to (partially) mitigate it.
@@
Naturally, reducing the marking intensity would attenuate the model’s memorization on the marked samples, and increase their prediction loss. But we find that the prediction loss on the member samples are still much lower than that on non-members, and thus MembershipTracker can overcome it via the set-based MI process (similar to Section 5.2.1), and maintain a 0% FPR@100% TPR [^4] However, there are certain scenarios where the reduced data-marking intensity would lead to lower auditing performance, and an evaluation for this is in Appendix C.1.2..
Figure 8. Unsupervised OOD detection. How different data-marking intensities affect the detection performance. Left: Samples marked with higher marking intensity are prone to be detected (0.8085 AUC). Right: Reducing the marking intensity can effectively reduce the AUC to only 0.6333.
Supervised OOD detection. Next, we consider a more knowledgeable defender who is aware of the potential data marking strategy. We use the MCHAD method to train a supervised OOD detector (Kirchheim et al., 2022a), and we consider two scenarios below.
@@
Figure 9. Supervised OOD detection. Blue line assumes a defender with perfect knowledge of the type of outlier features adopted by the users for data marking. Orange line considers a realistic defender with knowledge of the possible types of outlier features, but the exact one followed by the users to mark their data is unpredictable.
@@
Specifically, we train the detector on samples marked with random color stripes, and evaluate it on samples marked with the features from an OOD (TinyImageNet) dataset (this is an effective alternative strategy we study in Appendix C.1.3). The result is in the orange line in Fig. 9, and the detection performance degrades considerably under such a realistic setting (with only 0.66 AUC).
@@
Spurious features detection is another related method. We evaluate a leading technique by Neuhaus et al. (Neuhaus et al., 2023) based on neural PCA components. But we find that it also has limited success in detecting the marked samples (0.6181 AUC).
5.2.7. Output perturbation
MembershipTracker leverages the model’s outputs to perform MI auditing, and we consider two strategies to perturb the model’s outputs: (1) injecting Gaussian noise to the outputs; and (2) returning only the top-1 prediction label (Choquette-Choo et al., 2021; Huang et al., 2024). However, we find that MembershipTracker still maintains high auditing success even with the perturbed outputs.
Figure 10. Output noise injection increases the prediction loss of the member samples (from the leftmost to the middle figure), but their loss are still lower than many non-member samples’. Thus, MembershipTracker can leverage the average loss of the samples by each user to achieve a 0% FPR@ 100% TPR (the rightmost figure).
6. Evaluation on ImageNet Training
This section evaluates MembershipTracker under the full-sized ImageNet-1k training (over 1.28 million training samples). We train a ResNet50 (He et al., 2016) and Swin-Transformer (Liu et al., 2021), and they have 75.09% and 80.29% accuracy respectively (both with <1% accuracy drop). We consider the challenging setting of supporting multiple users in different classes (1,000 target users in total). Each user contributes 125 samples to the training set, which is $<$0.01% in proportion.
We find that even under the large-scale training setting, the marked samples can still be strongly memorized by the models, which enables MembershipTracker to achieve 0% FPR@100% TPR.